# On the Characteristics of the Conjugate Function Enabling Effective Dual Decomposition Methods

### v1

@@ Appendix B Lipschitzian Properties of $f^{*}$ and $g$
We derive Lipschitzian properties of the conjugate function $f^{*}$ of $f$ [cf.  (2)] and the dual function $g$ of problem (2) under Assumption 1. Since we consider strictly convex functions, our results are more general than the existing Lipschitzian properties for strongly convex objective functions [42, Prop. 12.60]. We first outline some results that are useful when asserting the intended results.
@@
Suppose Assumption 1 holds. Then the conjugate function $f^{*}$ of $f$ is lsc, proper, and strictly convex with $\texttt{dom}~{}f^{*}={\rm I\!R}^{n}$.
@@
The result follows directly from Theorem 11.1 of [42], together with that $f^{*}(\bm{\nu})$ is finite for all $\bm{\nu}\in{\rm I\!R}^{n}$. ∎
@@
Since Assumption 1 holds, the function $f$ is almost strictly convex, in the sense that $f$ is strictly convex in $\texttt{rint}~{}\mathcal{Y}$, cf. Definition 5. Thus, Theorem 11.13 of [42] guarantees that $f^{*}$ is almost differentiable on $\texttt{int}~{}(\texttt{dom}~{}f^{*})={\rm I\!R}^{n}$ [cf. Lemma 2], which in turn guarantees that $f^{*}$ is differentiable on ${\rm I\!R}^{n}$. ∎
@@
Proposition 11.3 of [42], together with Assumption 1 and Lemma 3 ensures that
@@
TABLE I: Summary of Results in Figure 5LABEL:sub@subfig:linear-regression-8 and Figure 8LABEL:sub@subfig:linear-regression-4
Figure Method $k^{\star}$ $k_{0}$ $b^{\star}$
Figure 5LABEL:sub@subfig:linear-regression-8 FGOR +{+}+ SD $115$ $21$ $105+6580b$
SD $421$ $NA$ $29470b$
FGOR $+$ Ref [35] $61$ $5$ $25+3920b$
Ref [35] $183$ $NA$ $12810b$
Figure 8LABEL:sub@subfig:linear-regression-4 FGOR $+$ SD $451$ $49$ $343+30954b$
SD $>1000$ $a$ $NA$ $>77000b$
FGOR $+$ Ref [35] $57$ $25$ $175+2464b$
Ref [35] $\gg 1000$ $b$ $NA$ $\gg 77000b$
  • $a$
    For $k=1000$ iterations, $g(\bm{\lambda}^{(k)})-g(\bm{\lambda}^{\star})=5.4471\times 10^{-4}$.
  • $b$
    For $k=1000$ iterations, $g(\bm{\lambda}^{(k)})-g(\bm{\lambda}^{\star})=0.524$.
@@
Define $\texttt{lip}~{}F(\bm{\nu})$, the Lipschitz modulus of a single-valued mapping $F:{\rm I\!R}^{m}\rightarrow{\rm I\!R}^{n}$ at $\bm{\nu}$ [42, Def. 9.1]:
@@
The central step in the proof is to show that $\forall{\bm{\nu}\in{\rm I\!R}^{n}}$, $\texttt{lip}~{}\nabla f^{*}(\bm{\nu})\leq L$, which in turn implies the desired result from [42, Theorem 9.2]. To this end, for each vector $\mathbf{u}=[u_{1}\ u_{2}\ \ldots\ u_{n}]^{\mbox{\scriptsize T}}\in{\rm I\!R}^{n}$, define the function $(\mathbf{u}\nabla f^{*}):{\rm I\!R}^{n}\rightarrow{\rm I\!R}$ by
@@
$\sup_{\|\mathbf{u}\|_{2}=1}\texttt{lip}~{}(\mathbf{u}\nabla f^{*})(\bm{\nu})$ $=\sup_{\|\mathbf{u}\|_{2}=1}\texttt{lip}\left(\sum_{i=1}^{n}u_{i}\frac{\partial f^{*}(\bm{\nu})}{\partial\nu_{i}}\right)$ (45)
@@
The equality (45) follows from (44), (46) follows from [42, Exerc. 9.8(b)], (47) follows from [42, Exerc. 9.8(a)], and (48) follows from Lemma 4 and (43) since the scalar-valued function $\partial f^{*}(\bm{\nu})/\partial\nu_{i}$ is Lipschitz continuous with $\max_{\mathbf{y}\in\mathcal{Y}}\|\mathbf{y}\|$, (49) follows from trivial rearrangements of terms, and (50) follows directly from the fact that $\sup~{}\|\mathbf{u}\|_{1}$ is achieved when $\mathbf{u}=[1/\sqrt{n}\ \ldots\ 1/\sqrt{n}]^{\mbox{\scriptsize T}}$. Finally, we note that $\texttt{lip}~{}\nabla f^{*}(\bm{\nu})=\sup_{\|\mathbf{u}\|_{2}=1}\texttt{lip}~{}(\mathbf{u}\nabla f^{*})(\bm{\nu})$ to conclude the result, cf. [42, Exerc. 9.9]. ∎
It is worth noting that the Lipschitzian properties claimed in Proposition 2 do not rely on any strong convexity properties of $f$. However, if such properties are imposed on $f$, Lipschitzian properties of $\nabla f^{*}$ directly follow from the results pertaining to the dualization of strong convexity, cf.  [42, Theorem 12.60]. The result is summarized in the following Remark.
@@
Suppose Assumption 1 holds. Then the dual function $g$ is differentiable on ${\rm I\!R}^{m}$. The gradient $\nabla g$ of $g$ is Lipschitz continuous with the constant $G=||A||^{2}_{2}L$.
@@
Appendix C Proof of Remark 4
Recall that $\mathbf{y}^{(k)}$ is the minimizer of the Lagrangian associated to problem (2) with Lagrange multiplier $\bm{\lambda}^{(k)}$. Then from [42, Prop. 11.3], we have

### v3

@@ Appendix F Lipschitzian Properties of $f^{*}$ and $g$
We derive Lipschitzian properties of the conjugate function $f^{*}$ of $f$ [cf.  (2)] and the dual function $g$ of problem (2) under Assumption 1. Since we consider strictly convex functions, our results are more general than the existing Lipschitzian properties for strongly convex objective functions [Rockafellar-98, Prop. 12.60]. We first outline some results that are useful when asserting the intended results.
@@
Suppose Assumption 1 holds. Then the conjugate function $f^{*}$ of $f$ is lower semicontinuous (lsc), proper, and strictly convex with $\texttt{dom}~f^{*}={\rm I\!R}^{n}$.
@@
The result follows directly from Theorem 11.1 of [Rockafellar-98], together with that $f^{*}(\bm{\nu})$ is finite for all $\bm{\nu}\in{\rm I\!R}^{n}$. ∎
@@
Since Assumption 1 holds, the function $f$ is almost strictly convex, in the sense that $f$ is strictly convex in $\texttt{rint}~\mathcal{Y}$, cf. Definition 5. Thus, Theorem 11.13 of [Rockafellar-98] guarantees that $f^{*}$ is almost differentiable on $\texttt{int}~(\texttt{dom}~f^{*})={\rm I\!R}^{n}$ [cf. Lemma 3], which in turn guarantees that $f^{*}$ is differentiable on ${\rm I\!R}^{n}$. ∎
@@
Proposition 11.3 of [Rockafellar-98], together with Assumption 1 and Lemma 4 ensures that
@@
Define $\texttt{lip}~F(\bm{\nu})$, the Lipschitz modulus of a single-valued mapping $F:{\rm I\!R}^{m}\rightarrow{\rm I\!R}^{n}$ at $\bm{\nu}$ [Rockafellar-98, Def. 9.1]:
@@
The central step in the proof is to show that $\forall{\bm{\nu}\in{\rm I\!R}^{n}}$, $\texttt{lip}~\nabla f^{*}(\bm{\nu})\leq L$, which in turn implies the desired result from [Rockafellar-98, Theorem 9.2]. To this end, for each vector $\mathbf{u}=[u_{1}\ u_{2}\ \ldots\ u_{n}]^{\mbox{\scriptsize T}}\in{\rm I\!R}^{n}$, define the function $(\mathbf{u}\nabla f^{*}):{\rm I\!R}^{n}\rightarrow{\rm I\!R}$ by
@@
$\hskip-8.53581pt\sup_{\|\mathbf{u}\|_{2}=1}\texttt{lip}~(\mathbf{u}\nabla f^{*})(\bm{\nu})$ $=\sup_{\|\mathbf{u}\|_{2}=1}\texttt{lip}\left(\sum_{i=1}^{n}u_{i}\frac{\partial f^{*}(\bm{\nu})}{\partial\nu_{i}}\right)$ (55)
@@
The equality (55) follows from (54), (56) follows from [Rockafellar-98, Exerc. 9.8(b)], (57) follows from [Rockafellar-98, Exerc. 9.8(a)], and (58) follows from Lemma 5 and (53) since the scalar-valued function $\partial f^{*}(\bm{\nu})/\partial\nu_{i}$ is Lipschitz continuous with $\max_{\mathbf{y}\in\mathcal{Y}}\|\mathbf{y}\|$, (59) follows from trivial rearrangements of terms, and (60) follows directly from the fact that $\sup~\|\mathbf{u}\|_{1}$ is achieved when $\mathbf{u}=[1/\sqrt{n}\ \ldots\ 1/\sqrt{n}]^{\mbox{\scriptsize T}}$. Finally, we note that $\texttt{lip}~\nabla f^{*}(\bm{\nu})=\sup_{\|\mathbf{u}\|_{2}=1}\texttt{lip}~(\mathbf{u}\nabla f^{*})(\bm{\nu})$ to conclude the result, cf. [Rockafellar-98, Exerc. 9.9]. ∎
It is worth noting that the Lipschitzian properties claimed in Proposition 3 do not rely on any strong convexity properties of $f$. However, if such properties are imposed on $f$, Lipschitzian properties of $\nabla f^{*}$ directly follow from the results pertaining to the dualization of strong convexity, cf.  [Rockafellar-98, Theorem 12.60]. The result is summarized in the following Remark.
@@
[…]